OpenStudy (anonymous):

I think its no solution. Please help :) Solve the given system, using the substitution method. y = 3x – 4 9x – 3y = 14 A. There is no solution. B. (12, 14) C. (14, 12) D. There are an infinite number of solutions.

OpenStudy (anonymous):

There's a variety of ways to substitute, but the first equation already defines y explicitly as a function of x. In other words, y is already by itself. So, we'll plug it into the second equation: 9x - 3 [ 3x - 4 ] = 14 9x - 9x - 12 = 14 -12 = 14. Woah, that's not right. -12 is not equal to 14. So, we've reached a false statement. That means that there are no solutions between these two lines.

OpenStudy (anonymous):

Oh woops. That should be a +12 instead of a -12. I forgot that the -3 multiplies with the -4 to yield positive 12. Either way, the solution is still inconsistent since 12 is not equal to 14.